# Mediation analysis in longitudinal intervention studies with an ordinal treatment-dependent confounder

### v1

@@ 5 Discussion
We addressed causal mediation in the case of a longitudinal mediator and a time-to-event outcome in a randomised lifestyle intervention study, where part of the lifestyle changes constituted a treatment-dependent confounder. Such situations may arise in interventional studies which target at reducing the risk for an adverse health outcome through lifestyle changes. When the interest lies in the treatment effect mediated through changes in a biological risk factor, such as body weight in the current study, other lifestyle changes will act as treatment-dependent confounders complicating the identification of mediational effects. The methods described here can be applied to address such situations in order to elucidate the mechanisms by which treatments operate on health outcomes.
@@
Previously, Tchetgen Tchetgen and VanderWeele (2014) showed that if the effect of treatment on a binary-valued treatment-dependent confounder is monotonic, mediational effects can be identified from empirical data. We here extended their approach to an ordinal trichotomous case, where a single sensitivity parameter needs to be specified for identification of the mediational effects. The sensitivity parameter determines the joint probability distribution of the two counterfactuals of the treatment-dependent confounder and identifies the component of the total effect controlling how the effect due to the additive interaction of the mediator and the treatment-dependent confounder is divided into the direct and indirect effects. Importantly, the consistency of the monotonicity assumption with observed data can be assessed based on the restrictions it enforces on the observed distribution of the treatment-dependent confounder.
@@
Other options to deal with treatment-dependent confounding have been introduced. Identification can be retained if the counterfactuals of the treatment-dependent confounder are assumed to be independent or if one is a deterministic function of the other (Robins and Richardson, 2010), or if there is no additive interaction between the treatment and the treatment-dependent confounder (Tchetgen Tchetgen and VanderWeele, 2014). Another option would be to switch the estimation target from natural direct and indirect effects to their interventional analogues, which are identifiable even in the presence of treatment-dependent confounders but do not, in general, have the same interpretation as the natural effects (VanderWeele, 2014b; Nguyen et al., 2022; Miles, 2023).
@@
We treated the longitudinal mediator (BMI) as a functional entity, i.e. a smooth function describing the underlying trajectory which the mediator follows and of which observations endowed with stochastic deviations were made over time. The contribution of the stochastic deviations to the association between the mediator and the time-to-event outcome was assumed to be negligible, hence promoting the underlying function itself as the effective mediator. An implicit assumption was that the structure given for the trajectory function is flexible enough to capture the longitudinal process to a relevant extent. Defining the mediator as a functional entity avoided the practical difficulties dealing with a high-dimensional mediator and also the conceptual problems that might arise from having to control for post-treatment values of the mediator. Similar approaches treating longitudinally measured mediators as functional entities have been previously investigated in the mediation analysis literature (Lindquist, 2012; Coffmann et al., 2023; Zeng et al., 2023).
@@
We chose the restricted survival time as the time-to-event outcome to avoid issues arising from unmeasured heterogeneity under outcomes defined conditionally on previous survival (Aalen et al., 2015). As one might plausibly assume, for example, some unobserved genetic factor affecting T2D risk, conditioning on survival at any post-treatment time would likely induce a selection bias invalidating the analysis. In addition, as one third of the study participants did not develop T2D, a considerable tail proportion of the event time distribution remained unobserved, thus creating the risk of misspecifying the parametric shape of the tail distribution. Therefore, we chose the restriction time as 15 years since the baseline to represent a clinically meaningful time period providing reasonable number of follow-up visits and T2D events.
@@
We used a joint modelling framework to estimate the parametric models. The association structure was induced by including the latent mediator trajectory in the survival submodel and also specifying a shared random effects structure (Gould et al., 2015; Papageorgiou et al., 2019). As the mediator was considered a latent quantity, individual-specific mediator trajectories needed to be estimated and the uncertainty regarding the estimation appropriately propagated to the estimates of the causal effects. This propagation of uncertainty was straightforwardly handled in the Bayesian set-up of computations.
@@
From the causal perspective, shared random effects between the longitudinal and survival submodels correspond to latent confounding between the mediator and the outcome. The joint modelling framework can thus be employed to account for such unmeasured confounding (Zheng and Liu, 2021). We here used the random intercept of the mediator trajectory as a shared random effect to reflect a latent property of the individual that might contribute to the level of the mediator trajectory as well as the time-to-event outcome. In our empirical application this might be, for example, an unmeasured metabolism-related genetic factor.
@@
The empirical application motivating this study aimed at decomposing the treatment effect on the restricted mean T2D-free time into indirect (mediated through weight reduction) and direct (all other mechanisms) effects. With the three-year legacy parameterisation, the estimated indirect effect translated to one year of additional time without T2D over the 15 years after the start of follow-up, with a $95\%$ credible interval ranging from 1.5 months to two years. The estimated total effect was 1.6 years with a $95\%$ credible interval from negative one month to positive three years. These results suggest that the early weight reduction may indeed constitute a major mechanism by which a lifestyle intervention affects long-term survival from T2D, accounting for two thirds of the total effect as calculated crudely from the point estimates. With the current change parameterisation, however, the indirect effect was estimated to be considerably smaller (0.31 years, $95\%$ CI 0.080.08-0.08- 0.080.750.750.750.75 years), accounting, based on the point estimates, roughly $20\%$ of the total effect (1.6 years, $95\%$ CI 0.310.310.310.313.13.13.13.1).
@@
Asuming that the treatment effects cannot be negative had little effect on the estimates under the current change parameterisation, and the proportion mediated was estimated to be $23\%$ with the $95\%$ credible interval ranging from $2\%$ to $64\%$. With the three-year legacy parameterisation, however, $23\%$ of the numerical posterior samples were discarded and the direct effect estimate increased and the indirect effect estimate decreased compared with the unrestricted case. The decrease of the indirect effect estimate after removing samples with negative effects was due to the fact that the two mediational effects had a strong negative correlation and, since the posterior distribution of direct effect extended to the negative side, a large number of samples with a very large negative direct effect, and consequently a very large positive indirect effect, were not admissible.
@@
The choice of the legacy parameterisation was based on previous literature showing the persistence of the intervention effects on T2D risk long after the discontinuation of the active treatment and loss of the acquired group differences in the clinical risk factors (Wilding, 2014). We used the cumulative change at three years as a time-constant predictor in the survival submodel starting from the treatment onset. We justify this by noting that the individual trajectories are determined by the observed baseline covariates and the random effects and can thus be interpreted to exist at the baseline, even though they may be learned only by observing the trajectory unfold over time. In this context, the weight reduction over the first three years was interpreted as a surrogate for some biological process which reacts quickly to the initiation of the lifestyle intervention and the induced behavioural changes. This underlying biological process was then assumed to manifest itself through the subsequent weight reduction and be the true causal mechanism associating weight reduction and T2D risk.
@@
Direct and indirect effects were identified up to a sensitivity parameter controlling the probability ($p_{11}$) that the two counterfactual levels of the trichotomous treatment-dependent confounder both belong to the middle category (corresponding to an absent treatment effect). The estimated causal effects remained practically unaffected by the chosen level of the sensitivity parameter. This was likely due to the fact that the constraint imposed by the monotonicity assumption on $p_{11}$ (see eq. (3)) was very tight. Moreover, increasing the probability of an absent treatment effect necessarily also increases the probability ($p_{02}$) of a large treatment effect and decreases the probabilities for intermediate treatment effects ($p_{01}$ and $p_{12}$). If the treatment-dependent confounder has a monotonic effect on the mediator and the outcome, this may be seen as a kind of a self-regulating property, since we must be always be up- or downweighting the highest and lowest effects together.
@@
Our empirical analysis has some potential limitations. First, the lifestyle score acting as the treatment-dependent confounder was formed rather crudely by summarising variables describing the targeted lifestyle changes over the first three follow-up visits. As the treatment-dependent confounder is an important part of the causal mechanism, any inaccuracy in its measurement may distort the resulting causal effect estimates. Second, the generalisability of the results are limited. The DPS inclusion criteria accepted volunteering individuals who were already overweight, had developed an impaired glucose tolerance and were between 40 to 65 years of age at the screening visit but had not yet been diagnosed with T2D. The DPS cohort thus represents a selected population, that is, individuals with a high risk of type 2 diabetes but who have managed to avoid the disease up to a relatively old age.
@@
In conclusion, we investigated causal mediation in longitudinal intervention studies with a time-to-event outcome in the presence of an ordinal treatment-dependent confounder. Foremost, we showed that assuming monotonicity of the treatment effect on a trichotomous ordinal treatment-dependent confounder, the direct and indirect effects can be identified up to a scalar sensitivity parameter. The time-to-event outcome was defined as a restricted survival time to avoid issues pertaining to measures conditioning on previous survival. We treated the longitudinal mediator as a functional entity to overcome problems with a high-dimensional mediator and employed a joint modelling framework to control for possible unobserved confounding between the mediator and the outcome. The methodology was applied to decompose the effect of a lifestyle intervention on restricted T2D-free time into parts mediated through weight reduction (indirect effect) and other mechanisms (direct effect). We found some evidence in favour of the existence of a clinically significant indirect effect through weight reduction, however, the magnitude of the estimated indirect effect depended considerably on the assumed effective form of the mediator. Choosing the weight change over the first three years as the mediator, the indirect effect accounted for a large fraction of the total effect, while taking the current weight change since baseline as the mediator, the direct effect was much larger than the indirect one.
Funding
MV was financially supported by the EXACTUS programme in the University of Turku Graduate School (UTUGS).

### v2

@@ 5 Discussion
We addressed causal mediation in the case of a longitudinal mediator and a time-to-event outcome in a randomised lifestyle intervention study, where some of the lifestyle changes acted as treatment-dependent confounders. Such situations may arise in interventional studies aimed at reducing the risk for an adverse health outcome through lifestyle changes. When the interest lies in the treatment effect mediated through changes in a biological risk factor, such as body weight in the current study, other lifestyle changes will act as treatment-dependent confounders, complicating the identification of mediational effects. The methods described here can be applied to address such situations in order to elucidate mechanisms by which treatments affect health outcomes.
@@
Previously, Tchetgen Tchetgen and VanderWeele (2014) showed that if the effect of treatment on a binary-valued treatment-dependent confounder is monotonic, mediational effects can be identified from empirical data. We here extended their approach to an ordinal trichotomous case, where a single sensitivity parameter (for each baseline covariate stratum) needs to be specified to identify the mediational effects. The sensitivity parameter determines the joint probability distribution of the two counterfactuals, $(U_{a^{*}},U_{a})$, of the treatment-dependent confounder and identifies the component of the total effect that controls how the effect due to the additive interaction of the mediator and the treatment-dependent confounder is divided into the direct and indirect effects. Our approach is related to partial identification, where bounds for the effect estimates can be determined by optimising the equations (2) with respect to the unidentified joint distribution without any structural assumptions on the joint distribution of $(U_{a^{*}},U_{a})$ (Miles et al., 2017). The monotonicity assumption may be seen as a structural assumption that sets some of the joint probabilities to zero. This also implies a necessary condition on the marginals of the joint distribution of $(U_{a^{*}},U_{a})$, allowing empirical assessment of the tenability of the monotonicity assumption.
@@
Other approaches to deal with treatment-dependent confounding have been introduced. Identification can be retained if the counterfactuals of the treatment-dependent confounder are assumed to be independent or if one is a deterministic function of the other (Robins and Richardson, 2010), or if there is no additive interaction between the treatment and the treatment-dependent confounder (Tchetgen Tchetgen and VanderWeele, 2014). Another option would be to switch the estimation target from natural direct and indirect effects to their interventional analogues, which are identifiable even in the presence of treatment-dependent confounders but do not, in general, share the same interpretation as the natural effects (VanderWeele, 2014b; Nguyen et al., 2022; Miles, 2023).
@@
We treated the longitudinal mediator (BMI) as a functional entity, i.e. a smooth function describing the underlying trajectory which the mediator follows and of which observations endowed with stochastic deviations were made over time. The contribution of the stochastic deviations to the association between the mediator and the time-to-event outcome was assumed to be negligible, hence promoting the underlying function itself as the effective mediator. An implicit assumption was that the structure given for the trajectory function is flexible enough to capture the longitudinal process to a relevant extent. By defining the mediator as a functional entity, we avoided the practical difficulties dealing with a high-dimensional mediator, as well as the conceptual challenges that might arise from having to control for post-treatment values of the mediator. Similar approaches treating longitudinally measured mediators as functional entities have been previously investigated in the mediation analysis literature (Lindquist, 2012; Coffmann et al., 2023; Zeng et al., 2023).
@@
We chose the restricted survival time as the time-to-event outcome to avoid issues arising from unmeasured heterogeneity under outcomes defined conditionally on previous survival (Aalen et al., 2015). For instance, if there were an unobserved genetic factor influencing T2D risk, conditioning on survival at any post-treatment time would induce a selection bias, potentially invalidating the analysis. In addition, as one third of the study participants did not develop T2D, a substantial tail proportion of the event time distribution remained unobserved. This may introduce the risk of misspecifying the parametric shape of the tail distribution. To address this, we chose the restriction time as 15 years since the baseline to represent a clinically meaningful time period providing reasonable number of follow-up visits and T2D events.
@@
We used a joint modelling framework to estimate the parametric models. The association structure was induced by incorporating the latent mediator trajectory in the survival submodel and also specifying a shared random effects structure (Gould et al., 2015; Papageorgiou et al., 2019). As the mediator was considered a latent quantity, individual-specific mediator trajectories needed to be estimated and the uncertainty regarding the estimation appropriately propagated to the estimates of the causal effects. This propagation of uncertainty was straightforwardly handled in the Bayesian set-up of computations.
@@
From the causal perspective, shared random effects between the longitudinal and survival submodels represent latent confounding between the mediator and the outcome. The joint modelling framework can thus be employed to account for such unmeasured confounding (Zheng and Liu, 2021). We here used the random intercept of the mediator trajectory as a shared random effect to reflect a latent property of the individual that may influence both the mediator trajectory and the time-to-event outcome. In our empirical application this might be, for example, an unmeasured metabolism-related genetic factor.
@@
The primary goal of our empirical application was to decompose the treatment effect on the restricted mean T2D-free time into indirect (mediated through weight reduction) and direct (all other mechanisms) effects. Using the three-year legacy parameterisation, the estimated indirect effect translated to one year of additional time without T2D over the 15 years after the start of follow-up, with a $95\%$ credible interval ranging from 1.5 months to two years. The estimated total effect was 1.6 years with a $95\%$ credible interval from negative one month to positive three years. These results suggest that the early weight reduction may indeed constitute a major mechanism through which a lifestyle intervention affects long-term T2D-free survival, accounting for two thirds of the total effect as calculated crudely from the point estimates. However, using the current change parameterisation, the indirect effect was estimated to be considerably smaller (0.31 years, $95\%$ CI 0.080.08-0.08- 0.080.750.750.750.75 years), accounting, based on the point estimates, roughly $20\%$ of the total effect (1.6 years, $95\%$ CI 0.310.310.310.313.13.13.13.1). Sensitivity analyses relaxing the monotonicity assumption yelded largly consistent results.
@@
Asuming that the treatment effects cannot be negative had little impact on the estimates under the current change parameterisation, and the proportion mediated was estimated to be $23\%$ with the $95\%$ credible interval ranging from $2\%$ to $64\%$. With the three-year legacy parameterisation, however, $23\%$ of the numerical posterior samples were discarded and the direct effect estimate increased and the indirect effect estimate decreased compared with the unrestricted case. The decrease of the indirect effect estimate after removing samples with negative effects was due to the fact that the two mediational effects had a strong negative correlation. Since the posterior distribution of direct effect extended far to the negative side, a large number of samples with a very large negative direct effect, and consequently a very large positive indirect effect, were not admissible.
@@
The choice of the legacy parameterisation was motivated by previous literature demonstrating the persistence of the intervention effects on T2D risk long after the discontinuation of the active treatment and diminishing of the acquired group differences in the clinical risk factors (Wilding, 2014). We used the cumulative change at three years as a time-constant predictor in the survival submodel starting from the onset of treatment. We justify this by noting that the individual trajectories are determined by the observed baseline covariates and the random effects and can thus be interpreted to exist at the baseline, even though they may be learned only by observing the trajectory unfold over time. In this context, the weight reduction during the first three years was interpreted as a surrogate for some biological process that responds quickly to the initiation of the lifestyle intervention and the induced behavioural changes. This underlying biological process was then assumed to manifest through the subsequent weight reduction and be the true causal mechanism linking weight reduction to a decrease in T2D risk.
@@
The direct and indirect effects were identified up to sensitivity parameters controlling the probabilities ($p_{11}$) that the two counterfactual levels of the trichotomous treatment-dependent confounder in each stratum both belong to the middle category (corresponding to an absent treatment effect). The estimated causal effects remained largely unaffected by the sensitivity parameter. This was likely due to the fact that the constraint imposed by the monotonicity assumption on $p_{11}$ (see eq. (3)) was very stringent. Furthermore, increasing the probability of an absent treatment effect necessarily also increases the probability ($p_{02}$) of a large treatment effect, simultaneously decreasing the probabilities for intermediate treatment effects ($p_{01}$ and $p_{12}$). If the treatment-dependent confounder has a monotonic effect on the mediator and the outcome, this may be seen as a self-regulating property, since the highest and lowest effects must always be up- or downweighted together.
@@
Our empirical analysis has some potential limitations. First, the lifestyle score acting as the treatment-dependent confounder was constructed somewhat crudely by summarising variables describing the targeted lifestyle changes over the first three follow-up visits. As the treatment-dependent confounder is an important part of the causal mechanism, any inaccuracy in its measurement may distort the resulting causal effect estimates. Second, the generalisability of the results is limited. The DPS inclusion criteria selected volunteers who were already overweight, had developed impaired glucose tolerance and were between 40 to 65 years of age at the screening visit but had not yet been diagnosed with T2D. As such, the DPS cohort represents a selected population, that is, individuals at a high risk of T2D who had managed to avoid the disease until a relatively old age.
@@
In conclusion, we investigated causal mediation in longitudinal intervention studies with a time-to-event outcome in the presence of an ordinal treatment-dependent confounder. Foremost, we showed that assuming monotonicity of the treatment effect on a trichotomous ordinal treatment-dependent confounder, the direct and indirect effects can be identified up to stratum-specific scalar sensitivity parameters. The time-to-event outcome was defined as a restricted survival time to avoid issues pertaining to measures conditioning on prior survival. To overcome challenges with a high-dimensional mediator, we treated the longitudinal mediator as a functional entity and employed a joint modelling framework to control for possible unobserved confounding between the mediator and the outcome. The methodology was applied to decompose the effect of a lifestyle intervention on restricted T2D-free time into an indirect effect through weight reduction and a direct effect involving other mechanisms. We found some evidence suggesting the existence of a clinically significant indirect effect through weight reduction, however, the magnitude of the estimated indirect effect depended considerably on the assumed effective form of the mediator. When using the weight change over the first three years as the mediator, the indirect effect accounted for a large fraction of the total effect. Conversely, when considering the current weight change since baseline as the mediator, the direct effect was substantially larger than the indirect one. The results remained similar in sensitivity analyses relaxing the monotonicity, indicating robustness to violations of this assumption.